# Finite presentability of outer automorphism groups of large-type Artin groups

### v1

@@ Abstract
Given a connected large-type Artin group $A_{\Gamma}$, we introduce a deformation space $\mathcal{D}$. If $\Gamma$ is triangle-free, or has all labels at least 6, we show that this space is canonical, in that it depends only on the isomorphism type of $A_{\Gamma}$, and admits an $\mathrm{Out}(A_{\Gamma})$-action. Using this action we conclude that $\mathrm{Out}(A_{\Gamma})$ is finitely presentable. We emphasise that our proof can handle cases where $\Gamma$ has separating vertices, which were previously problematic.
In fact, our proof works for all connected large-type Artin groups satisfying the technical condition of having rigid chunks. We conjecture that all connected large-type Artin groups have rigid chunks, and therefore finitely presented outer automorphism groups.
@@ 1 Introduction
In the present work, we show that many large-type Artin groups have finitely presentable outer automorphism groups. Our main result is the following.
@@
(Theorem 4.1) Suppose $A$ is a connected large-type Artin group with rigid chunks. Then $\mathrm{Out}(A)$ is finitely presentable.
The property of having rigid chunks is a technical condition, and we postone the precise definition to Definitions 3.2 and 3.3. Essentially, this condition means subgraphs of defining graphs for $A$ with no separating edges or vertices (the so called chunks) are isomorphism invariant. We conjecture this technical condition is always satisfied in the connected large-type setting.
@@
Of course, if the conjecture is true, it immediately follows from our main theorem that $\mathrm{Out}(A)$ is finitely presentable for any connected large-type Artin group $A$.
@@
(Corollary 4.2) Suppose $A_{\Gamma}$ is a connected large-type Artin group, and $\Gamma$ either is triangle-free, or has every label at least 6. Then $\mathrm{Out}(A_{\Gamma})$ is finitely presentable.
In general, automorphism groups can be complicated even (or perhaps especially) for seemingly simple groups. For example, the (outer) automorphism group of $\langle x,t\mid tx^{2}t^{-1}=x^{4}\rangle$ is not even finitely generated [Lev07].
For a connected large-type Artin group $A$ with rigid chunks, we study $\mathrm{Out}(A)$ via its action on a (finite dimensional deformation retract of a) contractible simplicial complex called a deformation space, a notion originally introduced by Forester [For01]. To our knowledge, this is the first time deformation spaces have been considered in the context of Artin groups. Before providing further details of our strategy, we will discuss the history of studying outer automorphisms of Artin groups, motivating the adoption of this technology.
@@ 1.1 Historical Context
Free groups are some of the simplest Artin groups. The study of the automorphism groups of free groups predates even the definition of Artin groups, and continues to be an active area. In [Nie24] it was shown that the outer automorphism groups of free groups are finitely presentable. Later, Culler and Vogtmann introduced the so called Outer Space, a contractible topological space parameterising free $F_{n}$-actions on trees, with an $\mathrm{Out}(F_{n})$-action [VC86]. This provided an alternative proof of finite presentability. The outer space has been a fundamental tool in the study of $\mathrm{Out}(F_{n})$ since its introduction.
More recently, Bregman-Charney-Vogtmann have introduced a generalisation of outer space for right-angled Artin groups (RAAGs) [BCV23]. This space parameterises actions on complexes generalising the well known Salvetti complexes. Like the classical outer space, the existence of this space demonstrates that RAAGs have finitely presented outer automorphism groups (previously known by work of Day [Day09]), and promises answers to further questions such as the outer automorphism groups’ finite subgroups.
@@
One is that of the spherical Artin groups, that is those where the associated Coxeter group quotient is finite, and affine Artin groups. The earliest result in this direction was finding the automorphism groups of braid groups [DG81], and the outer automorphism groups of spherical dihedral Artin groups have also been known for over 20 years [GHMR00]. Recently the (outer) automorphism groups for Artin groups of type $D_{n}$, $\tilde{A_{n}}$ and $B_{n}$ have been determined [CP24] [PS24b][PS24a]. We will not focus on this direction since, in some sense, spherical Artin groups are as far from large as possible.
The other direction is the two-dimensional Artin groups, of which the large-type Artin groups are a subclass. Crisp initiated the study of automorphism groups of two-dimensional Artin groups, finding a finite generating set for the isomorphism groupoid of connected large-type triangle-free (that is to say, $A_{\Gamma}$ of connected large-type where the graph $\Gamma$ is triangle-free) Artin groups [Cri05]. It quickly follows that the automorphism groups are finitely generated. Blufstein, Vaskou and Martin recently proved the same result for XXXL-type Artin groups [BMV24]. Crisp also showed that, in the case that $\Gamma$ (once again connected, large-type and triangle-free) also has no separating vertices, $\mathrm{Out}(\Gamma)$ is virtually abelian. Recently, An and Cho found explicit presentations in this case [AC22].
@@
The Artin groups in the previous paragraph are rigid in the sense that if $A_{\Gamma}\cong A_{\Gamma^{\prime}}$ where $\Gamma$ admits a twistless heirarchy terminating in twistless stars, then $\Gamma\cong\Gamma^{\prime}$. Morally, this is what limits the number of automorphisms and results in $\mathrm{Out}(A_{\Gamma})$ being finite. Notably, these defining graphs have no separating vertices or edges. The Artin groups addressed by An and Cho may have separating edges (but not vertices), and it is possible that $A_{\Gamma}\cong A_{\Gamma^{\prime}}$ where $\Gamma$ and $\Gamma^{\prime}$ are not isomorphic but “twisted” over one of these separating edges. These twists commute, which justifies the outer automorphism groups being virtually abelian. As such, these groups can still be thought of as towards the rigid end of the spectrum of large-type Artin groups.
@@ 1.2 Our methods
To prove Theorem 1.1, we show that the action of $\mathrm{Out}(A)$ on $\mathcal{F}$ is co-compact, and that the simplex stabilisers are finitely presented, with the more involved part being co-compactness. Finite presentability of $\mathrm{Out}(A)$ follows immediately.
The canonical deformation space provided in Theorem 1.4 is a promising tool for the study of outer automorphism groups of large-type Artin groups beyond the question of finite presentability. For example, the outer space for a free group has the Nielsen realisation property: that is, any finite subgroup of $\mathrm{Out}(F_{n})$ fixes a point in the associated outer space [Zim81]. This property is named in analogy with the corresponding fact about mapping class groups acting on Teichmüller space [Ker83]. The same result holds for finite subgroups of the so called untwisted subgroup of $\mathrm{Out}(A)$ where $A$ is a general RAAG [BCV24]. An immediate consequence of our Theorem 1.4 is a weak form of Nielsen realisation for large-type Artin groups.
@@ Acknowledgements
The author thanks their supervisor, Laura Ciobanu, for her support and reading of the manuscript. This project would not have been completed without many helpful conversations, thank you to Naomi Andrew, Ruth Charney, Ric Wade, Giovanni Satori and Alexandre Martin.
@@ 2 Preliminaries
This section is divided into two subesections covering known results we will use. Section 2.1 is about Artin groups, and 2.2 is about deformation spaces.
@@ 2.1 Artin Groups
A defining graph $\Gamma$ is a finite simplicial graph where every edge has a label in $\mathbb{Z}_{\geq 2}$, which we denote $m_{ab}$ where $a,b$ are the endpoints of the edge.
@@
Given a defining graph $\Gamma$, the corresponding Artin group $A_{\Gamma}$ is defined to be the group presented by
@@
Now, we will say $A$ is an abstract Artin group if $A\cong A_{\Gamma}$ for some defining graph $\Gamma$. Each such defining graph $\Gamma$ will be called a defining graph for $A$.
@@
Take $\Gamma_{1}$ a defining graph with all labels at least 3, then it is known if $A_{\Gamma_{1}}\cong A_{\Gamma_{2}}$, $\Gamma_{2}$ also has all labels at least 3 [MV24]. Moreover, if $\Gamma_{1}$ has all labels at least 6 (resp. is triangle free, is connected) then the same is true of $\Gamma_{2}$ [Vas23b]. As such, in each of the previous definitions one could replace the word “some” with “any”.
@@
This motivates our interest in dihedral Artin groups. For $m\in\mathbb{N}_{\geq 3}$ we will write
$DA_{m}=\langle s,t\mid\ \underbrace{sts\cdots}_{m\text{ terms}}=\underbrace{tst\cdots}_{m\text{ terms}}\rangle.$
We will sometimes call spherical dihedral Artin groups odd or even based on whether $m$ is odd or even.
@@
[God07, Corollary 4.12] Suppose $A_{\Gamma}$ is a large-type Artin group and $A_{\Lambda}\leq A_{\Gamma}$ is a spherical dihedral parabolic subgroup. Suppose that $g\in A_{\Gamma}$ and $gA_{\Lambda}g^{-1}=A_{\Lambda^{\prime}}$. Then $\Lambda=\Lambda^{\prime}$ and $g\in A_{\Lambda}$.
@@
[Par97, Corollary 4.2] Suppose $A_{\Gamma}$ is a large-type Artin group, and $s,t$ are standard generators. Suppose that $g\in A_{\Gamma}$ and $gA_{t}g^{-1}\leq A_{s}$. Then in $\Gamma$ there exists a path of odd edges of $\Gamma$ through vertices $s=s_{0},s_{1}\dots s_{n}=t$, and $g\in Z(A_{s})\Delta_{s_{0},s_{1}}\Delta_{s_{1},s_{2}}\dots\Delta_{s_{n-1},s_{n}}$.
@@ 2.2 Deformation Spaces
The appropriate equivalence relation for $G$-trees is $G$-equivariant isometry. If $(T_{1},\Omega_{1})$ and $(T_{2},\Omega_{2})$ are $G$-trees, an isometry $f:T_{1}\rightarrow T_{2}$ is said to be $G$-equivariant if for all $g\in G$, $f\circ\Omega_{1}(g)=\Omega_{2}(g)\circ f$.
Notice if $\varphi_{h}$ is conjugation by $h$, then $(T,\Omega)$ and $(T,\Omega\circ\varphi_{h})$ are equivariantly isometric via the isometry $\Omega(h)$. This observation will allow an action of $\mathrm{Aut}(G)$ on a deformation space to descend to an action of $\mathrm{Out}(G)$.
@@
The following is a useful characterisation of when two trees are in the same deformation space. In the following, for a $G$-tree $T$, we write $l_{T}:G\rightarrow\mathbb{R}$ for the corresponding length function, where $l_{T}(g)=\inf\{d(x,gx)\mid x\in T\}$.
@@
[For01, Theorem 1.1] Let $G$ be a group and $X$ and $Y$ be cocompact $G$-trees. Then the following are equivalent:
@@
Slide moves are important due to the following theorem.
@@
One attractive feature of deformation spaces is that they are contractible. For us, this will mean that a group $G$ acting on $\mathcal{D}$ is the fundamental group of the complex of groups arising from the quotient. We note that contractibility of deformation spaces is a general fact, and holds without further assumptions such as $T/G$ being a tree.
@@
Our final tool will be a subcomplex of $\mathcal{D}$ constructed by Guirardel and Levitt, which is a deformation retract of $\mathcal{D}$ and therefore still contractible, and is finite dimensional.

### v2

# Type VF for outer automorphism groups of large-type Artin groups

@@ Abstract
Given a connected large-type Artin group $A_{\Gamma}$, we introduce a deformation space $\mathcal{D}$. If $\Gamma$ is triangle-free, or has all labels at least 6, we show that this space is canonical, in that it depends only on the isomorphism type of $A_{\Gamma}$, and admits an $\mathrm{Out}(A_{\Gamma})$-action. Using this action we conclude that $\mathrm{Out}(A_{\Gamma})$ is of type VF, which implies $\mathrm{Out}(A_{\Gamma})$ finitely presentable. We emphasise that our proof can handle cases where $\Gamma$ has separating vertices, which were previously problematic.
In fact, our proof works for all connected large-type Artin groups satisfying the technical condition of having rigid chunks. We conjecture that all connected large-type Artin groups have rigid chunks, and therefore outer automorphism groups of type VF.
@@ 1 Introduction
In the present work, we show that the outer automorphism groups of many large-type Artin groups have type VF, meaning there is a finite index subgroup with a finite classifying space (see Definition 2.24 and Definition 2.25). This is the strongest possible finiteness condition, in particular implying finite generatability and finite presentability.
@@
(Theorem 4.1) Suppose $A$ is a connected large-type Artin group with rigid chunks. Then $\mathrm{Out}(A)$ is of type VF.
The property of having rigid chunks is a technical condition, and we postpone the precise definition to Definitions 3.2 and 3.3. Essentially, this condition means subgraphs of defining graphs for $A$ with no separating edges or vertices (the so called chunks) are isomorphism invariant. We conjecture this technical condition is always satisfied in the connected large-type setting.
@@
Of course, if this conjecture is true, it immediately follows from our main theorem that $\mathrm{Out}(A)$ is of type VF for any connected large-type Artin group $A$. We remark that Crisp conjectured a generating set for the automorphism group of a connected large-type Artin group [Cri05], and that if Crisp’s conjecture is true, then so is Conjecture 1.2.
@@
(Corollary 4.2) Suppose $A_{\Gamma}$ is a connected large-type Artin group, and $\Gamma$ either is triangle-free, or has every label at least 6. Then $\mathrm{Out}(A_{\Gamma})$ is of type VF.
In any case of the previous corollary where $\Gamma$ has a separating vertex, finite presentability of $\mathrm{Out}(A_{\Gamma})$ is new. If $\Gamma$ is furthermore not triangle-free, even finite generatability is new.
We remark that in general, automorphism groups can be complicated even (or perhaps especially) for seemingly simple groups. For example, the (outer) automorphism group of $\langle x,t\mid tx^{2}t^{-1}=x^{4}\rangle$ is not even finitely generated [CL83].
For a connected large-type Artin group $A$ with rigid chunks, we study $\mathrm{Out}(A)$ via its action on a (finite dimensional deformation retract of a) contractible simplicial complex called a deformation space, a notion originally introduced by Forester [For01]. To our knowledge, this is the first time deformation spaces have been considered in the context of outer automorphisms of Artin groups. Before providing further details of our strategy, we will discuss the history of studying outer automorphisms of Artin groups, motivating the adoption of this technology.
@@ 1.1 Historical Context
Free groups are some of the simplest Artin groups. The study of the automorphism groups of free groups predates even the definition of Artin groups, and continues to be an active area. In [Nie24] it was shown that the outer automorphism groups of free groups are finitely presented. Later, Culler and Vogtmann introduced the so called Outer Space, a contractible topological space parameterising free $F_{n}$-actions on trees, with an $\mathrm{Out}(F_{n})$-action [VC86]. In the same paper, they used this space (and its cocompact spine) to show $\mathrm{Out}(F_{n})$ is of type VF. The outer space has been a fundamental tool in the study of $\mathrm{Out}(F_{n})$ since its introduction.
Charney, Stambaugh and Vogtmann have now introduced a generalisation of outer space for so called untwisted automorphisms of right-angled Artin groups (RAAGs) [CSV17]. This space is parameterised by actions on complexes generalising the well known Salvetti complexes. Like the classical outer space, the existence of this space, and a cocompact spine, demonstrates that, for each RAAG $A_{\Gamma}$, the untwisted subgroup $U(A_{\Gamma})\leq\mathrm{Out}(A_{\Gamma})$ is of type VF. More recently, Bregman, Charney and Vogtmann introduced a full outer space for RAAGs, although there is no cocompact spine. Nonetheless, Day and Wade showed, via different methods, $\mathrm{Out}(A_{\Gamma})$ is of type VF for each RAAG $A_{\Gamma}$ [DW19].
@@
One is that of the spherical Artin groups, that is those where the associated Coxeter group quotient is finite, and affine Artin groups. The earliest result in this direction was finding the automorphism groups of braid groups [DG81]. Around 20 years ago, the outer automorphism groups of spherical dihedral Artin groups [GHMR00], and Artin groups of type $B_{n}$, $\tilde{A_{n}}$ and $\tilde{C_{n}}$ were also calculated [CC05] (see also [PS24b][PS24a]). Recently the (outer) automorphism groups for Artin groups of type $D_{n}$ were determined [CP24]. We will not focus on this direction since, in some sense, spherical Artin groups are as far from large as possible, but remark in all cases listed in this paragraph the outer automorphism groups are virtually abelian (often they are finite).
The other direction is the two-dimensional Artin groups, of which the large-type Artin groups are a subclass. Crisp initiated the study of automorphism groups of two-dimensional Artin groups, finding a finite generating set for the isomorphism groupoid of connected large-type triangle-free (that is to say, $A_{\Gamma}$ of connected large-type where the graph $\Gamma$ is triangle-free) Artin groups [Cri05]. It quickly follows that the automorphism groups are finitely generated. Blufstein, Vaskou and Martin recently proved the same result for XXXL-type Artin groups without separating vertices in $\Gamma$ [BMV24]. Crisp also showed that, in the case that $\Gamma$ (once again connected, large-type and triangle-free) also has no separating vertices, $\mathrm{Out}(\Gamma)$ is virtually abelian. Recently, An and Cho found explicit presentations in this case [AC22].
@@
The Artin groups in the previous paragraph are rigid in the sense that if $A_{\Gamma}\cong A_{\Gamma^{\prime}}$ where $A_{\Gamma}$ is large-type and admits a twistless heirarchy terminating in twistless stars, then $\Gamma\cong\Gamma^{\prime}$. Morally, this is what limits the number of automorphisms and results in $\mathrm{Out}(A_{\Gamma})$ being finite. Notably, these defining graphs have no separating vertices or edges. The Artin groups addressed by An and Cho may have separating edges (but not vertices), and it is possible that $A_{\Gamma}\cong A_{\Gamma^{\prime}}$ where $\Gamma$ and $\Gamma^{\prime}$ are not isomorphic but “twisted” over one of these separating edges. These twists commute, which justifies the outer automorphism groups being virtually abelian. As such, these groups can still be thought of as towards the rigid end of the spectrum of large-type Artin groups.
@@ 1.2 Our methods
To prove Theorem 1.1 the first, and most involved, step is to show that the action of $\mathrm{Out}(A)$ on $\mathcal{F}$ is co-compact. Then, we pass to a finite index subgroup $\mathrm{Out}_{0,+}(A)\leq\mathrm{Out}(A)$, which is well defined only because of the rigid chunks. Finally we show that $\mathrm{Out}_{0,+}(A)$ acts with type F simplex stabilisers. Type VF for $\mathrm{Out}(A)$ follows immediately.
The canonical deformation space provided in Theorem 1.4 is a promising tool for the study of outer automorphism groups of large-type Artin groups beyond studying finiteness properties. For example, the outer space for a free group has the Nielsen realisation property: that is, any finite subgroup of $\mathrm{Out}(F_{n})$ fixes a point in the associated outer space [Zim81]. This property is named in analogy with the corresponding fact about mapping class groups acting on Teichmüller space [Ker83]. The same result holds for finite subgroups of the untwisted subgroup of $U(A)\leq\mathrm{Out}(A)$ where $A$ is a general RAAG [BCV24]. An immediate consequence of our Theorem 1.4 is a weak form of Nielsen realisation for large-type Artin groups.
@@ Acknowledgements
The author thanks their supervisor, Laura Ciobanu, for her support and reading of the manuscript. This project would not have been completed without many helpful conversations, thank you to Naomi Andrew, Ruth Charney, Giovanni Sartori and Alexandre Martin. Thank you to Ignat Soroko and Gilbert Levitt for their comments on the first version, and thanks especially to Ric Wade, who suggested the approach to strengthen the main result from finite presentability to type VF.
@@ 2 Preliminaries
This section is divided into three subesections covering known results we will use. Section 2.1 is about Artin groups, and Section 2.2 is about deformation spaces, and Section 2.3 is about the finiteness property VF.
@@ 2.1 Artin groups
A defining graph $\Gamma$ is a finite simplicial graph where every edge $\{a,b\}$ has a label in $m_{ab}\in\mathbb{Z}_{\geq 2}$.
@@
Given a defining graph $\Gamma$, the corresponding Artin group $A_{\Gamma}$ is defined to be the group with the following presentation:
@@
Now, we will say $A$ is an abstract Artin group if $A\cong A_{\Gamma}$ for some defining graph $\Gamma$. Each such defining graph $\Gamma$ will be called a defining graph for $A$. Note that we may have $\Gamma_{1}\neq\Gamma_{2}$ such that $A_{\Gamma_{1}}\cong A_{\Gamma_{2}}$.
@@
Take $\Gamma_{1}$ a defining graph with all labels at least 3, then it is known if $A_{\Gamma_{1}}\cong A_{\Gamma_{2}}$, $\Gamma_{2}$ also has all labels at least 3 [MV24]. Moreover, if $\Gamma_{1}$ also has all labels at least 6 (resp. is triangle free, is connected) then the same is true of $\Gamma_{2}$ [Vas23b]. As such, in each of the previous definitions one could replace the word “some” with “any”.
@@
This motivates our interest in dihedral Artin groups. For $m\in\mathbb{N}_{\geq 3}$ we will write,
$DA_{m}=\langle s,t\mid\ \underbrace{sts\cdots}_{m\text{ terms}}=\underbrace{tst\cdots}_{m\text{ terms}}\rangle,$
for the Artin group defined by a single edge labelled by $m$. We will sometimes call spherical dihedral Artin groups odd or even based on whether $m$ is odd or even.
@@
Lemma 2.9.
Suppose $H\leq\mathrm{Out}(DA_{2m})$ fixes $\langle b\rangle$ up to conjugation. Then $H$ is finite.
Proof.
We apply the isomorphism,
$\langle a,b\mid(ab)^{n}=(ba)^{n}\rangle\cong\langle x,t\mid x^{n}=tx^{n}t^{-1}\rangle,$
given by $b\mapsto t$ and $a\mapsto xt^{-1}$.
By [GHMR00, Theorem D], $\mathrm{Out}(DA_{n})$ is generated by inner automorphisms and the following:
$\alpha$ $:x\mapsto x^{-1},t\mapsto t$
$\beta$ $:x\mapsto x,t\mapsto t^{-1}$
$\gamma$ $:x\mapsto x,t\mapsto tx,$
and in particular any outer automorphism class has a unique representative $\alpha^{i}\beta^{j}\gamma^{k}$ where $i,j\in\{0,1\}$ and $k\in\mathbb{Z}$.
Note if $\phi=\alpha^{i}\beta^{j}\gamma^{k}$ fixes $\langle t\rangle$ (the image of $b$) up to conjugation, then the induced map on the abelianisation (which one easily checks is $\langle\bar{x},\bar{t}\mid[\bar{x},\bar{t}]\rangle$) fixes $\langle\bar{t}\rangle$. In particular, this means $k=0$. The result follows. ∎
@@
[God07, Corollary 4.12] Suppose $A_{\Gamma}$ is a large-type Artin group and $A_{\Lambda}\leq A_{\Gamma}$ is a parabolic subgroup with at least 2. Suppose that $g\in A_{\Gamma}$ and $gA_{\Lambda}g^{-1}=A_{\Lambda^{\prime}}$. Then $\Lambda=\Lambda^{\prime}$ and $g\in A_{\Lambda}$.
@@
[Par97, Corollary 4.2] Suppose $A_{\Gamma}$ is a large-type Artin group, and $s,t$ are standard generators. Suppose that $g\in A_{\Gamma}$ and $gA_{t}g^{-1}\leq A_{s}$. Then in $\Gamma$ there exists a path of odd edges of $\Gamma$ through vertices $s=s_{0},s_{1}\dots s_{n}=t$, and $g\in Z_{A_{\Gamma}}(A_{s})\Delta_{s_{0},s_{1}}\Delta_{s_{1},s_{2}}\dots\Delta_{s_{n-1},s_{n}}$.
@@ 2.2 Deformation spaces
The appropriate equivalence relation for $G$-trees is $G$-equivariant isometry.
Definition 2.14.
If $(T_{1},\Omega_{1})$ and $(T_{2},\Omega_{2})$ are $G$-trees, an isometry $f:T_{1}\rightarrow T_{2}$ is said to be $G$-equivariant if for all $g\in G$, $f\circ\Omega_{1}(g)=\Omega_{2}(g)\circ f$.
Notice if $\iota_{h}$ is conjugation by $h$, then $(T,\Omega)$ and $(T,\Omega\circ\iota_{h})$ are equivariantly isometric via the isometry $\Omega(h)$. This observation will allow an action of $\mathrm{Aut}(G)$ on a deformation space to descend to an action of $\mathrm{Out}(G)$.
@@
The following is a useful characterisation of when two trees are in the same deformation space. In the following, for a $G$-tree $T$, we will say $H\leq G$ is an elliptic subgroup if it acts on $T$ with a fixed point.
@@
[For01, Theorem 1.1] Let $G$ be a group and $T_{1}$ and $T_{2}$ be co-compact $G$-trees. Then the following are equivalent:
@@
Slide moves are important due to the following theorem. In the following, the running assumption that the quotient of each $G$-tree by $G$ is a tree is used.
@@
One attractive feature of deformation spaces is that they are contractible. For us, this will mean that a group $G$ acting on $\mathcal{D}$ is the fundamental group of the complex of groups arising from the quotient.
@@
Our final tool will be a subcomplex of $\mathcal{D}$ constructed by Guirardel and Levitt, which is a deformation retract of $\mathcal{D}$ and therefore still contractible, and is finite dimensional. Again, here the running assumption that $T/G$ is a tree is used.
@@
2.3 Type VF
Being type F is the strongest of the finiteness conditions. It in particular implies finite generatability and finite presentability.
Definition 2.24.
A group $G$ is of type F if there is a finite Eilenberg-MacLane space; that is, a CW-complex $X$ with finitely many cells such that $\pi_{1}(X)\cong G$ and each higher homotopy group is trivial.
We note that being type F is closed under finite direct products (by taking the direct product of the Eilenberg-MacLane spaces) and taking finite index subgroups (by taking a finite cover).
A group with torison can never be of type F, so the strongest condition we can ask for a group with torsion is that it has a (necessarily torsion free) finite index subgroup of type F. This motivates the definition of type VF.
Definition 2.25.
We say $G$ is of type VF if it has a finite index subgroup of type F.
The following theorem, applied with $X$ being a deformation space, will be our main tool for showing outer automorphism groups of certain Artin groups are of type VF. In the following rigid means any cell fixed setwise by the action of some $g\in G$ is fixed pointwise by $g$.
Theorem 2.26.
[Geo07, Theorem 7.3.4] Let $X$ be a contractible rigid $G$-CW complex whose quotient $X/G$ is finite. Let the stabilizer of each cell of $X$ have type F. Then G has type F.